x+2y-z=6
-x+y+3z=5
x+2y+2z=9
Im not sure what its called exactly what were working on, but you have to solve for x,y, and z in all of the equations.

Answer by KMST(5328)

It's a system of linear equations.
You have 3 pieces of information (the 3 equations), and 3 variables to be found.
An easier system would look like a staircase, like this:
x=1
-x+y=2
x+2y-z=6
That would be easier because the first equation gives you the value for one of the variable; plugging that value in the second equation will lead you to find the value for another variable, and substituting the values found in the third equation will allow you to find the remaining variable.
You can transform your system by combining equations. We pick one of the equations that we will leave untouched, and we'll make combinations of the chosen one with the other two.
-x+y+3z=5 will be kept
Adding up
x+2y-z=6 and
-x+y+3z=5 yields 3y+2z=11
Adding up
x+2y+2z=9 and
-x+y+3z=5 yields 3y+5z=14
We will continue to work with the equivalent system below
-x+y+3z=5 (which will be kept)
3y+2z=11
3y+5z=14
With the two last equations we will pick one to be kept untouched and make a combination of the two to replace the other
Keeping 3y+5z=14,
we multiply 3y+2z=11 times (-1) to get
-3y-2z=-11 and add
3y+5z=14 yielding 3z=3 (which is the same as z=1).
Now we have
3z=3 (or z=1)
3y+5z=14
-x+y+3z=5
Can you solve that one?